37. A continuous flow water heater (geyser) has an electrical power rating 2 kW and efficienty of conversion of electrical power into heat-80%. If water is flowing through the device at the rate of 100 cc/sec, and the inlet temperature is 10°C, the outlet temperature will be (1) 12.2°C (3) 20°C (2) 13.8°C (4) 16.5°C
Answer #1

Here ,

Power , P = 2000 W

power converted to heat = 0.80 * P

power converted to heat = 0.80 * 2000

power converted to heat = 1600 W

mass of water entering per s = 100 gm = 0.10 Kg

let the outlet temperature is T

heat = m * S * delta T

1600 = 0.10 * 4186 * (T - 10)

solving for T

T = 13.8 degree C

the outlet temperature is 2) 13.8 degree C